A 100N weight rest on a 30 degrees inclined plane. Neglecting friction, how much pull must one exert to bring the weight up the plane


1
Expert's answer
2022-01-18T09:35:06-0500

"\text{Component of gravity acting along an inclined plane:}"

"F_g = mg\sin\alpha"

"\alpha= 30\degree"

"mg = 100N"

"F_{min} = F_g"

"F_{min} = 100\sin30\degree= 50N"


"\text{Answer:} 50N"